OpenStudy (anonymous):

A sphere with a diameter of 12 has a volume of Xπ cubic units. A square has an area of X sq. units. What is the perimeter of the square? please show me the steps.

OpenStudy (dumbcow):

volume sphere = 4/3*pi*r^3 r = half diameter = 6 X = 4/3*6^3 = 4/3*216 = 288 if square has sides of s, then s^2 = X=288 s = sqrt(288) = 16.97 perimeter = 4s = 67.88

OpenStudy (anonymous):

answer is 48 sqrt 2

OpenStudy (dumbcow):

right i cheated and used calculator :) s = sqrt(288) = sqrt(144)sqrt(2) = 12sqrt2 perimeter = 4s = 48sqrt(2)

OpenStudy (anonymous):

Thanks for helping me.!

OpenStudy (dumbcow):

:)